# The fate of a single impurity in the Bose-Hubbard model

### v1

@@ Abstract
Collectively, these observables delineate three dressing regimes driven solely by $U_{\mathrm{ib}}/t$: (i) a light-polaron regime (red), characterized by finite winding, a broad, weak depletion cloud extending over $R\!\gtrsim\!4-5$, and coherent quasiparticle motion; (ii) a heavy polaron regime (yellow), where the winding number has already collapsed ($\langle W_{\mathrm{imp}}^{2}\rangle\!\approx\!0$), yet $\Delta N(R)$ still displays a compact but finite-range deformation ($R\simeq 3$–4), signaling an immobile but still spatially dressed impurity; (iii) a saturated-bubble regime (blue), where all $\Delta N(R)$ curves beyond $R\simeq 2$ collapse onto a universal profile, indicating a fully localized defect with a spatially saturated depletion cloud. Unlike in the Mott background, this saturation is not quantized to 1-1, but deepens continuously with increasing $U_{\mathrm{ib}}/t$ while the bath remains globally superfluid. This progression from mobile light polarons to spatially saturated bubble represents a continuous, non-singular localization route driven purely by impurity-bath coupling strength inside a compressible superfluid bath.
@@
Figure 4: Compressibility-controlled collapse of polaronic dressing at a fixed repulsive impurity bath coupling $U_{\mathrm{ib}}/t=8.0$. (a) Ground-state energy $E_{\mathrm{p}}/t$, (b) effective-mass ratio $m^{*}/m_{0}$, and (c) quasiparticle residue $Z_{0}$ as functions of bath interaction $U_{\mathrm{b}}/t$. The shaded region marks the superfluid–MI transition of the bath ($U_{\mathrm{b}}/t\simeq 16.7$). In the compressible superfluid regime ($U_{\mathrm{b}}/t<16.7$), the impurity exhibits a weakly renormalized light-polaron behavior with finite dressing cloud. As the bath becomes increasingly less compressible, $m^{*}/m_{0}$ decreases while $Z_{0}$ rises, indicating a collapse of polaronic dressing. Once the bath enters the incompressible MI regime ($U_{\mathrm{b}}/t\gtrsim 16.7$), $m^{*}/m_{0}$ approaches unity and $Z_{0}\to 1$, confirming the emergence of a nearly-free.
@@
(i) Polaron renormalization under decreasing bath compressibility. To isolate the role of the bath, we fix the impurity–bath coupling at a moderate repulsive value ($U_{\mathrm{ib}}/t=8.0$) and vary $U_{\mathrm{b}}/t$ across the superfluid–MI transition. Figures 4(a–c) show that in the compressible superfluid regime ($U_{\mathrm{b}}/t<16.7$), the impurity remains a well-defined polaron with finite winding [Fig. 5(e)]. As the bath becomes less compressible, the many-body dressing weakens: the effective mass $m^{*}/m_{0}$ decreases from $\sim 1.5$ toward 1.2 while the quasiparticle residue $Z_{0}$ increases from $\sim 0.7$ to nearly 0.9. This trend reflects the fact that a less compressible bath is less able to mediate dressing, leading to a progressively lighter polaron, in contrast to the usual impurity-bath interaction-driven mass enhancement for a fixed bath interaction. Upon approaching the SF–MI boundary ($U_{\mathrm{b}}/t\simeq 16.7$), $E_{\mathrm{p}}/t$ exhibits a visible change in slope, while $m^{*}/m_{0}\to 1$ and $Z_{0}\to 1$, signaling the collapse of many-body dressing and the recovery of an almost free impurity. Once the bath enters the incompressible Mott regime ($U_{\mathrm{b}}/t\gtrsim 16.7$), extended density redistribution is prohibited, and the impurity can no longer sustain polaronic correlations. It reduces to a nearly free defect—a trivial particle with no spatial dressing and no quasiparticle renormalization.
@@
This compressibility-controlled breakdown of polaronic dressing is fundamentally different from the interaction-driven self-trapping observed at large $U_{\mathrm{ib}}/t$ in the superfluid regime: here, localization does not originate from strong coupling, but from the inability of the bath to support collective screening.
@@
Figure 5: Compressibility-controlled collapse of polaronic dressing at fixed repulsive impurity bath coupling $U_{\mathrm{ib}}/t=8.0$. (a) Cumulative bath-density depletion $\Delta N(R)$ and (b) impurity–centered correlator $C_{\mathrm{ib}}(R)$ for increasing bath interaction $U_{\mathrm{b}}/t$, color-coded from deep superfluid (dark red: $U_{\mathrm{b}}/t=8.0$) through intermediate (light red/orange: $U_{\mathrm{b}}/t=13.3,15.0$) to near the transition (light blue: $U_{\mathrm{b}}/t=17.0$) and deep Mott regime (dark blue: $U_{\mathrm{b}}/t=20.0,22.0,24.0$). (c,d) Peak position $R(\Delta N_{\min})$ and depth $\Delta N_{\min}(R)$, showing the loss of extended bath correlations. (e) Impurity winding number $\langle W_{\mathrm{imp}}^{2}\rangle$ increases sharply once the bath loses compressibility, signaling the loss of coherent propagation. (f) Bath superfluid density $\rho_{\mathrm{b}}$ confirms that the collapse is driven by the vanishing bath compressibility. Together these observables reveal a compressibility-controlled crossover from a mobile polaron to a nearly free defect, and finally to a localized particle-defect in the MI background.
(ii) Real-space collapse and defect formation. Figures 5(a,b) reveal how the real-space bath responses evolve under the same protocol. In the superfluid phase, $\Delta N(R)$ and $C_{\mathrm{ib}}(R)$ exhibit an extended depletion cloud characteristic of a coherent polaron. As the bath becomes less compressible, both the depth $\Delta N_{\mathrm{min}}(R)$ and spatial extent of the depletion $R(\Delta N_{\mathrm{min}})$ shrink continuously, as quantified in Figs. 5(c,d). This collapse coincides with a sharp incresase of the winding number (Fig. 5(e)), while the bath superfluid density vanishes (Fig. 5(f)), demonstrating that the breakdown is driven by the loss of compressibility. Inside the MI regime, the impurity first behaves as a nearly-free particle and then locks into a localized vacancy defect, characterized by quantized depletion $\Delta N(R)\to-1$. This represents the compressibility-controlled localization mechanism across the SF-MI transition, distinct from the continuous interaction-driven self-trapping in the superfluid.
@@
MI extension and outlook– In an incompressible MI bath, extended density redistribution is forbidden and neither mobile polarons nor bound clusters survive. Increasing $|U_{\mathrm{ib}}|/t$ then produces only short-range distortions, and quasiparticle signatures ($m^{*}/m_{0}$, $Z_{0}$, $\langle W_{\mathrm{imp}}^{2}\rangle$) revert toward their bare values. Deep in the Mott regime, strong coupling nucleates quantized particle or vacancy defects, signaled by a $\pm 1$ change of the bath occupation; a systematic analysis of this insulating branch is reported in Ref. [26].
@@
Several extensions are natural. Finite temperature should activate particle–hole excitations in the Mott background and induce a temperature-dependent delocalization crossover. Relaxing $t_{\mathrm{a}}\!=\!t_{\mathrm{b}}$ adds a powerful knob: $t_{\mathrm{a}}\!\ll\!t_{\mathrm{b}}$ approaches the static-defect limit, while $t_{\mathrm{a}}\!\gg\!t_{\mathrm{b}}$ realizes a fast impurity in a slow correlated bath. Finally, the two-impurity sector can directly address bipolaron formation and transport by mapping binding energy and effective mass across the $(U_{\mathrm{ib}}/t,\,U_{\mathrm{b}}/t)$ plane, with an optional direct impurity–impurity coupling to separate bare from bath-mediated forces—all within the same sign-problem-free QMC framework. More discussion of experimental realizations and detection protocols is provided in our companion paper [26].

### v2

@@ Abstract
Collectively, these observables delineate three regimes controlled solely by $U_{\mathrm{ib}}/t$: (i) a light-polaron regime (red) with finite winding and a broad, weak depletion cloud; (ii) a heavy-polaron regime (orange) where winding has collapsed but a compact finite-range deformation persists; and (iii) a saturated-bubble regime (blue) where the deformation profile is spatially saturated. This progression from mobile light polarons to a saturated bubble constitutes a continuous, non-singular self-trapping route driven purely by impurity–bath couplings within a compressible superfluid bath.
@@
Figure 4: Compressibility-controlled collapse of polaronic dressing at fixed repulsive impurity–bath coupling $U_{\mathrm{ib}}/t=8.0$. (a) Ground-state energy $E_{\mathrm{p}}/t$, (b) effective-mass ratio $m^{*}/m_{0}$, and (c) quasiparticle residue $Z_{0}$ as functions of bath interaction $U_{\mathrm{b}}/t$. The shaded region marks the bath SF–MI transition ($U_{\mathrm{b}}/t\simeq 16.7$). In the compressible SF ($U_{\mathrm{b}}/t<16.7$), the impurity forms a weakly renormalized light polaron with a finite dressing cloud. As the bath becomes less compressible, $m^{*}/m_{0}$ decreases while $Z_{0}$ increases, indicating a progressive collapse of polaronic dressing. Upon entering the incompressible MI ($U_{\mathrm{b}}/t\gtrsim 16.7$), $m^{*}/m_{0}\to 1$ and $Z_{0}\to 1$, signaling a nearly bare (almost free) impurity.
@@
(i) Polaron renormalization under decreasing bath compressibility. To isolate the role of the bath, we fix the impurity–bath coupling at a moderate repulsive value ($U_{\mathrm{ib}}/t=8.0$) and vary $U_{\mathrm{b}}/t$ across the SF–MI transition. Figures 4(a–c) show that in the compressible SF ($U_{\mathrm{b}}/t<16.7$) the impurity remains a well-defined polaron with finite winding [Fig. 5(e)]. As the bath becomes less compressible, the dressing weakens continuously: the effective-mass ratio $m^{*}/m_{0}$ decreases (from $\sim 1.5$ toward $\sim 1.2$), while the residue $Z_{0}$ increases (from $\sim 0.7$ toward $\sim 0.9$), indicating a progressively lighter and more weakly dressed polaron. Near the SF–MI boundary ($U_{\mathrm{b}}/t\simeq 16.7$), $E_{\mathrm{p}}/t$ exhibits a visible change in slope and the renormalization collapses, with $m^{*}/m_{0}\to 1$ and $Z_{0}\to 1$, signaling the recovery of an almost bare impurity. Inside the incompressible MI, extended density redistribution is suppressed and the bath can no longer sustain polaronic correlations, so the impurity behaves as a nearly free defect in the lattice background.
@@
This compressibility-controlled depolaronization is distinct from the interaction-driven self-trapping at large $|U_{\mathrm{ib}}|/t$ in the SF: here the dressing collapses because the bath loses compressibility, not because the impurity–bath coupling is strong.
@@
Figure 5: Compressibility-controlled collapse of real-space dressing and defect binding at fixed $U_{\mathrm{ib}}/t=8.0$. (a) Cumulative bath-density depletion $\Delta N(R)$ and (b) impurity-centered correlator $C_{\mathrm{ib}}(R)$ for increasing $U_{\mathrm{b}}/t$, color-coded from deep SF (dark red: $U_{\mathrm{b}}/t=8.0$) through intermediate (light red/orange: $U_{\mathrm{b}}/t=13.3,15.0$) to near the transition (light blue: $U_{\mathrm{b}}/t=17.0$) and deep MI (dark blue: $U_{\mathrm{b}}/t=20.0,22.0,24.0$). (c,d) Minimum position $R_{\min}$ and depth $\Delta N_{\min}$ extracted from $\Delta N(R)$, quantifying the collapse of extended bath response. (e) Impurity winding $\langle W_{\mathrm{imp}}^{2}\rangle$ increases across the SF–MI transition, indicating enhanced coherent impurity transport. (f) Bath superfluid density $\rho_{\mathrm{b}}$ vanishes at the transition. Together these observables reveal a compressibility-controlled crossover from a mobile polaron in the SF to an almost bare impurity near criticality, and, deep in the MI, binding to a quantized defect sector.
(ii) Real-space collapse and defect binding. Figures 5(a,b) show how the real-space bath response evolves under the same protocol. In the SF, $\Delta N(R)$ and $C_{\mathrm{ib}}(R)$ display an extended deformation cloud characteristic of a coherent polaron. As $U_{\mathrm{b}}/t$ increases and the bath becomes less compressible, both the depth $\Delta N_{\min}$ and the extent $R_{\min}$ shrink continuously [Figs. 5(c,d)], reflecting the collapse of long-wavelength density response. Concomitantly, the impurity winding $\langle W_{\mathrm{imp}}^{2}\rangle$ increases across the transition [Fig. 5(e)], while the bath superfluid density vanishes [Fig. 5(f)]. In the MI, the impurity is therefore nearly bare at intermediate couplings, whereas deeper in the MI it locks to a vacancy-type defect for repulsive $U_{\mathrm{ib}}$, signaled by a quantized depletion plateau with $\Delta N(R)$ approaching 1-1 (within our resolution).
This constitutes the compressibility-controlled route: dressing collapses as the bath becomes incompressible, and self-trapping in the MI proceeds via quantized defect binding, distinct from the continuous interaction-driven winding-collapse self-trapping within the superfluid.
@@
MI extension and outlook– In an incompressible MI bath, the gapped spectrum and vanishing compressibility strongly suppress long-range density rearrangements, quenching mobile polarons and disfavoring extended bound states. Increasing $|U_{\mathrm{ib}}|/t$ then produces only short-range distortions, and quasiparticle diagnostics ($m^{*}/m_{0}$, $Z_{0}$, and $\langle W_{\mathrm{imp}}^{2}\rangle$) revert toward their bare-defect limits. Deep in the Mott regime, sufficiently strong coupling nucleates quantized vacancy- or particle-type defects, manifested by discrete changes $\Delta N_{\mathrm{b}}=\pm 1$ in the total bath occupation; a systematic analysis of this insulating branch is reported in Ref. [27].
@@
Several extensions are natural. Finite temperature will activate particle–hole excitations in the Mott background and can induce a temperature-dependent delocalization (re-mobilization) crossover of the impurity–defect complex. Relaxing $t_{\mathrm{imp}}\!=\!t_{\mathrm{b}}$ adds a powerful knob: $t_{\mathrm{imp}}\!\ll\!t_{\mathrm{b}}$ approaches the static-defect limit, whereas $t_{\mathrm{imp}}\!\gg\!t_{\mathrm{b}}$ realizes a fast impurity in a slow, strongly correlated bath. Finally, the two-impurity sector can directly address bipolaron formation and transport by mapping the binding energy and effective mass across the $(U_{\mathrm{ib}}/t,\,U_{\mathrm{b}}/t)$ plane; an additional direct impurity–impurity interaction would further allow one to disentangle bare impurity-impurity interactions from bath-mediated interactions—all within the same sign-problem-free QMC framework. Further discussion of experimental realizations and detection protocols is provided in our companion paper [27].